A cash flow of $500,000 may be received by Cynthia Bennett Company in one year, two years, or three years, with probabilities of 30%, 50%, and 20%, respectively. The rate of interest on default risk-free investments is 8%. Required: What is the expected present value of Cynthia Bennett' cash flow? Round your answers to the nearest whole dollar. With details please.
Answer #1
Year 1 = $500,000 x 30% x PV(8%,1) $ 1,38,888.89
Year 2 = $500,000 x 50% x PV(8%,2) $ 2,14,334.71
Year 3 = $500,000 x 20% x PV(8%,3) $    79,383.22
Total Cash Flow $ 4,32,607
Year 1 = $500,000 x 30% x PV(8%,1) =PV(8%,1,0,-(500000*30%))
Year 2 = $500,000 x 50% x PV(8%,2) =PV(8%,2,0,-(500000*50%))
Year 3 = $500,000 x 20% x PV(8%,3) =PV(8%,3,0,-(500000*20%))
Total Cash Flow =SUM(B3:B5)